Parth (parthkohli):

How to derive \({\bf \vec{F}} = \dfrac{m(v_f - v_i)}{t}\) or \({\rm \vec{F}} \propto \dfrac{m(v_f - v_i)}{t}\), for that matter? Note, you do not have to use \({\bf \vec{F}} = m{\bf \vec{a}}\)

Parth (parthkohli):

@amistre64

OpenStudy (dls):

\[\LARGE F=\frac{dp}{dt}\] Can we use this? :P

Parth (parthkohli):

Good idea! ^_^ No.

OpenStudy (amistre64):

force is equal to change in momentum divided by change in time

OpenStudy (dls):

-_- @ParthKohli

Parth (parthkohli):

@amistre64 Isn't that just what DLS said...

OpenStudy (amistre64):

i was just seeing if i remembered how to interpret the question correctly :)

Parth (parthkohli):

\[F = \dfrac{\Delta p}{\Delta t} = \dfrac{\Delta mv}{\Delta t}\]Is the above true? How did you derive it?

OpenStudy (amistre64):

Ft = p = mv2-mv1

OpenStudy (amistre64):

\[\vec F\Delta t=\Delta \vec p = m(\vec v_f-\vec v_i)\]

OpenStudy (amistre64):

not sure what you mean by how you derive it.

Parth (parthkohli):

Does not seem so convincing. :-|

OpenStudy (amistre64):

do we have to first prove that momentum is a product of mass and velocity? i never know when i have to reinvent the wheel on these things

Parth (parthkohli):

No, not that one. lol, got it

OpenStudy (amistre64):

changing an objects momentum requires applying a force over a given interval of time, right?\[\vec F\Delta t=\Delta \vec p\]

Parth (parthkohli):

I'd rather say...\[\vec F \Delta t \propto \Delta \vec p\]

OpenStudy (amistre64):

that style went out with Descartes :)

Parth (parthkohli):

I feel like a newbie here!

Parth (parthkohli):

Can you give me a link to the derivation? I got how you mean \(\Delta \vec p =m(\vec v _f - \vec v_i)\) and there's only a little work left... but I didn't get how \(\vec F \Delta t = \Delta \vec p \).

OpenStudy (dls):

You are indirectly asking for the proof of F=ma which only Isaac Newton knows.

OpenStudy (dls):

I think @shubhamsrg can derive F=ma. Its a piece of cake for him.

Parth (parthkohli):

Of course, of course!

OpenStudy (amistre64):

an object remains at a constant velocity (change in velocity is acceleration) unless acted upon by a force. therefore in order to change an objects momentum, you would need to apply a constant force

OpenStudy (amistre64):

over a given time interval, the mass "acquires" a new velocity, and therefore a change in momentum. does that sound about right?